if cos c=( cos a – cos b) /(1-cosa cosb) then tan^2(c/2) tan^2(b/2) is

Grade:11

1 Answers

Nishant Vora IIT Patna
9 years ago
Hello student, Please find the answer to your question below

use the formula
cos\theta =\frac{1-tan^{2}\theta}{1+tan^{2}\theta}

Put in the original ewaution and solve it ….….….….….….…...